# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 04 Mar 2026 v2

@@ Abstract
We present the exact analytic solution of the Makeenko-Migdal loop equations, thereby solving planar QCD ($N_{c}\rightarrow\infty$) in the continuum limit. The solution is constructed by quantizing internal Majorana fermions (elves) on a rigid Hodge-dual minimal surface. We demonstrate that these worldsheet degrees of freedom provide the exact algebraic mechanism required to satisfy the unintegrated vector loop equations, with the Pauli principle enforcing the planar factorization. In the local limit, the theory reduces to a confining analytic twistor string. By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles. The simplest sector with just one pole yields the exact linear Regge spectrum $m^{2}=\frac{\pi\sigma}{4}(n+1/12)$, matching the experimental Regge trajectories for light mesons ($\pi,K$ trajectories with $n=4J$ and $\rho$ trajectory with n=4J2n=4J-2) within 95% confidence intervals. The empirically correct intercepts $\alpha_{\pi}(0)=-1/48$ and $\alpha_{\rho}(0)=+23/48$ emerge directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor pole monodromy, without any ad hoc assumptions of string vibrations. Ultimately, this exact geometric quantization explicitly realizes Witten’s Master Field as a critical classical trajectory in twistor space.
@@ 1 Introduction
The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the loop equations. At leading order in the $1/N_{c}$ expansion, the Makeenko-Migdal equations  [^1] Referred to hereafter as the MM equations. While brevity dictates this acronym, the foundational credit remains fully shared with my former student, Yuri Makeenko, whose hard work and physical intuition half a century ago made these derivations possible. exhibit the characteristic factorization properties of a string theory, yet a concrete, mathematically controlled continuum realization has remained elusive.
@@
The $S$-matrix poles are generated by non-compact flat valleys in the effective action, which lead to the divergence of the steepest descent integral. We demonstrate that these flat valleys are dynamically produced by the multi-valued monodromy of the holomorphic twistor fields; along these winding trajectories, higher-order functional corrections to the action are strictly absent. Furthermore, the discrete spectrum is classified by the topological number of twistor poles inside the unit circle. This topological sector is rigorously protected: the crossing of the unit circle by a twistor pole leads to a divergent action, creating an impenetrable topological barrier, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop $|z|=1$. By analytically continuing the action to the boundary and summing over the exact monodromy phases, we obtain an exact Bohr-Sommerfeld quantization condition. For the simplest sector with just one pole, this yields the spectrum:
$m^{2}=\frac{\pi\sigma}{4}\left(n+\frac{1}{12}\right)$ (1.1)
where $n=4J$ for the unnatural parity ($\pi,K$) trajectories and n=4J2n=4J-2 for the natural parity ($\rho$) trajectory. Remarkably, the intercepts 1/48,+23/48-1/48,+23/48 emerge directly from the 2D Liouville kinetic anomaly.
In a profound sense, the culminating result of this construction is the explicit geometric realization of Witten’s original "Master Field" hypothesis for large-$N_{c}$ QCD. While the Master Field does not manifest as a single classical gauge configuration in coordinate spacetime, we discover its exact representation in loop space. The true classical limit of Planar QCD is identified not as a matrix algebra, but as a rigid classical geometric object: a critical trajectory in twistor space, conceptually parallel to the spectral curves of Seiberg-Witten theory and accessible via modern geometric quantization.
@@
Section 13.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings. Section 15 discusses the semiclassical (WKB) limit and the emergence of linear Regge behavior.
@@ 5 Conformal metric as a local gauge parameter
$\nabla_{k}=\partial_{k}+\frac{\mathrm{i}}{2}\omega_{k}\sigma^{3},$ (5.2)
@@ 14.4 Stochastic Quantization and Complex Langevin Dynamics
15 The Linear Regge trajectories in the WKB limit
15.1 The Chiral Helicoid Saddle Point
In the semiclassical limit (large effective action), the path integral is dominated by the saddle point of the effective action. We look for a solution corresponding to a meson with large angular momentum $J$, which corresponds to a rotating momentum configuration.
Consider a momentum distribution rotating in the $(1,2)$ plane with frequency $\omega$:
$P_{1}(\tau)+\mathrm{i}P_{2}(\tau)=P_{0}e^{\mathrm{i}\omega\tau},\qquad P_{3}=P_{4}=0.$ (15.1)
This source term in the action (11.7) drives the spinor fields. We search for a solution that respects this rotational symmetry. The natural ansatz corresponds to the helicoid minimal surface described in Appendix B.4.
In the generic spinor formalism, with rescaling $w(z)=z^{-1/2}$, the conventional Helicoid is generated by $\mu(z)\propto(1,-z)^{T}$. However, as established in Appendix F, our geometric prescription requires enforcing the exact chiral (Hodge-dual) configuration. We freeze the right-handed twistor to a constant basis spinor:
$\lambda(z)=\begin{pmatrix}z^{-1}\\ 1\end{pmatrix},\qquad\mu(z)=\mathrm{i}\begin{pmatrix}1\\ 0\end{pmatrix}$ (15.2)
Substituting this into the null vector definition $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, we find that the spatial tangent vectors no longer identically vanish in the extra dimensions, like they did with conventional helicoid. Our Hodge dual minimal surface, with only one of two chiral Gauss maps (9.12), being active, and other being frozen, permanently breaks the 3D embedding constraint and allows the minimal surface to span the 4-dimensional space. Integrating this yields the logarithmic cut $f(z)\sim\log z$, which maps the complex plane to the infinite strip of the helicoid. The boundary value on the real axis $z=e^{\tau}$ gives a rotating coordinate vector, consistent with the rotating momentum source.
15.2 The Symmetrized Minkowski Helicoid
The Minkowski Helicoid is parameterized by time $\tau\in[0,T]$ and radius $r\in[-R,R]$, see [7]. The time $T$ tends to infinity in the end.
$X(\tau,r)=(\ r\cos\omega\tau,\ r\sin\omega\tau,0,\tau)$ (15.3)
The induced Minkowski metric for a generic string is $g_{ab}=\text{diag}(1-\omega^{2}r^{2},1)$.
Crucially, as derived in Appendix F, the conformal factor for the generic surface evaluates to $4\cos^{2}(\sigma_{M})$. In stark contrast, our chiral twistor evaluates to an induced Euclidean conformal factor $\exp\left(2\rho\right)\propto(1+e^{-2\sigma_{E}})$. Applying our physical prescription—taking the parity-even arithmetic mean of the Left-chiral (SD) and Right-chiral (ASD) configurations—upon double Wick rotation ($\sigma_{E}\to\mathrm{i}\sigma_{M}$), precisely isolates the real part:
$\text{Re}\left(1+e^{-2\mathrm{i}\sigma_{M}}\right)=1+\cos(2\sigma_{M})=2\cos^{2}(\sigma_{M})$ (15.4)
Thus, the effective proper area of the symmetrized chiral Minkowski Helicoid is exactly halved relative to the conventional generic string.
In Minkowski space we are not seeking to minimize the area in the geometric sense, as the area functional is not positive definite and the minimization problem is distinct. Instead, our criterion is algebraic: we require a solution to the Loop Equations for the Wilson loop. This demands the self-duality of the area derivative and the additivity of the area functional. We utilize the fact that the Euclidean Helicoid is a minimal surface for arbitrary angular velocity $\omega$. We obtain the Minkowski Helicoid via a double Wick rotation:
$x_{4}\to\mathrm{i}t,\qquad\omega_{E}\to-\mathrm{i}\omega.$ (15.5)
Under this transformation, the argument of the trigonometric functions in the coordinate parametrization becomes real ($\omega_{E}x_{4}\to\omega t$), mapping the Euclidean Helicoid to a real surface in Minkowski space. Crucially, the condition that $S[C]$ is a zero mode of the Loop Equation depends on algebraic identities (self-duality of the area derivative and additivity). By the Identity Theorem, these analytic relations, established for the Euclidean minimal surface, must hold for its analytic continuation into Minkowski space. Thus, the symmetrized Minkowski Helicoid remains a valid solution to the QCD Loop Equation, bypassing the ill-posed variational problem of minimizing area in a space with an indefinite metric.
Figure 14: The helicoid spanned by rotating $\bar{q}q$ pair connected by a rigid stick (string). This minimal surface bounded by a double helix, was discovered by Meusnier in 1785.
Another important point is that the Helicoid boundary loop (double helix) is infinitely smooth, lacking any cusps or self-intersections. This means that no UV divergent renormalization factors could potentially affect our computation of the meson Regge trajectory.
15.3 Semiclassical Motivation: Twisted Boundary Conditions
@@
Before solving the spectral equation, we justify the choice of the Helicoid geometry from the perspective of the semiclassical quark path integral.
Consider the Green’s function for a meson composed of scalar quarks with mass $m_{q}$. In the quenched approximation, the effective Minkowski action involves a sum over quark trajectories $C$, weighted by the quark mass term and the area of the confining minimal surface $\Sigma$ bounded by $C$, neglecting the perturbative factor444the perimeter renormalization terms in that factor are included in the physical quark mass.. This is relevant to our WKB approximation in the Wilson loop at large loops:
$\mathcal{Z}\sim\int\mathcal{D}C\,\exp\left(-\mathrm{i}m_{q}\oint_{C}ds-\mathrm{i}\sigma S_{\mathrm{Mink}}[\Sigma_{C}]\right)$ (15.6)
To extract the spectrum of states with angular momentum $J$, we impose twisted boundary conditions. We require the quark positions at time $t=T$ to be a rotation of their positions at $t=0$ by an angle $\Theta=\omega T$ in the rotation plane $xy$:
$(x+\mathrm{i}y)(T)=e^{\mathrm{i}\omega T}(x+\mathrm{i}y)(0)$ (15.7)
The angular momentum $J$ arises as the thermodynamic conjugate to the twist frequency $\omega$.
The ground state of this system is determined by the saddle point of the effective action. This requires the simultaneous minimization of the boundary length (quark term) and the surface area (gluon term).

### 12 Mar 2026 v3

@@ Abstract
We present the exact analytic solution of the Makeenko-Migdal loop equations, thereby solving planar QCD $(N_{c}\rightarrow\infty)$ in the continuum limit. The solution is constructed by quantizing internal Majorana fermions (elves) on a rigid Hodge-dual minimal surface. We demonstrate that these worldsheet degrees of freedom provide the exact algebraic mechanism required to satisfy the unintegrated vector loop equations, with the Pauli principle enforcing the planar factorization. In the local limit, the theory reduces to a confining analytic twistor string. By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles. The simplest sector with just one pole yields parametric Regge trajectories governed by trigonometric functions. While mathematically non-linear, these geometric trajectories are numerically remarkably close to linear, matching the experimental data for light mesons (the unshifted $\pi,K$ trajectories and the topologically shifted $\rho$ trajectory) within 95% confidence intervals. The empirically correct intercepts $\alpha_{\pi}(0)=-1/96$ and $\alpha_{\rho}(0)=+47/96$ emerge directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor branch point monodromy, without any ad hoc assumptions of string vibrations. Ultimately, this exact geometric quantization explicitly realizes the Master Field as a critical classical trajectory in twistor space.
@@ 1 Introduction
The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the MM loop equations [^1] The Makeenko-Migdal equations, referred to hereafter as the MM equations, exhibit the characteristic factorization properties of a string theory. While brevity dictates this acronym, the foundational credit remains fully shared with my former student, Yuri Makeenko, whose hard work and physical intuition half a century ago made these derivations possible.. A concrete, mathematically controlled continuum realization has remained elusive.
@@
The $S$-matrix poles are generated by non-compact flat valleys in the effective action, which lead to the divergence of the steepest descent integral. We demonstrate that these flat valleys are dynamically produced by the multi-valued monodromy of the holomorphic twistor fields; along these winding trajectories, higher-order functional corrections to the action are strictly absent. Furthermore, the discrete spectrum is classified by the topological number of twistor poles inside the unit circle. This topological sector is rigorously protected: the crossing of the unit circle by a twistor pole leads to a divergent action, creating an impenetrable topological barrier, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop $|z|=1$. By analytically continuing the action to the boundary and summing over the infinite number of windings, with $SU(2)$ rotated boundary conditions for twistors, after projecting on the fixed spin state we obtain an exact Bohr-Sommerfeld quantization condition. For the simplest sector with just one pole, this yields the unified parametric Regge trajectory:
$J=\frac{\sin^{4}(\pi a)}{2\pi^{2}a(1-a)}\frac{m^{2}}{\sigma}-\frac{q}{2}-\frac{a}{48};$ (1.1)
$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+24(2J+q)}$ (1.2)
The topological trace phase $q=0$ identifies the unshifted unnatural parity ($\pi$, $K$) trajectories, while q=1q=-1 rigidly shifts the natural parity ($\rho$) trajectory by $1/2$.
Remarkably, while mathematically non-linear, these trajectories are numerically very close to straight lines (see Figure 15). The empirically correct asymptotic intercepts $\alpha_{\pi}(0)=-1/96$ and $\alpha_{\rho}(0)=47/96$ emerge directly from the 2D Liouville kinetic anomaly. The asymptotic expansion of the Regge trajectory $J(s)$ at $s=m^{2}/\sigma>1$ reads:
$J(s)$ $=\frac{2s}{\pi^{2}}-\left(\frac{q}{2}+\frac{1}{96}\right)+\frac{B_{1}}{s}+\frac{B_{3}}{s^{3}}+\dots);$ (1.3)
$B_{1}$ $=\frac{\pi^{2}}{36864\left(\pi^{2}-2\right)}\approx 0.0000340208;$ (1.4)
$B_{3}$ $=\frac{\pi^{6}\left(48-24\pi^{2}+5\pi^{4}\right)}{32614907904\left(\pi^{2}-2\right)^{4}}\approx 2.29*10^{-9}$ (1.5)
We interpret this construction as a geometric realization of the large-$N_{c}$ Master Field [54]. While the Master Field does not manifest as a single classical gauge configuration in coordinate spacetime, it admits an explicit formulation in loop space. The true classical limit of planar QCD is thus identified not as a matrix algebra, but as a rigid classical geometric object: a critical trajectory in twistor space, conceptually parallel to the spectral curves of Seiberg-Witten theory[45]. This geometric structure is accessible via exact WKB analysis [50, 12], where twistor monodromies dynamically generate flat valleys within the Lefschetz thimbles [55, 3].
@@
Section 13.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings.
@@ 5 Conformal metric as a local gauge parameter
$\nabla_{k}=\partial_{k}+\frac{\mathrm{i}}{2}\omega_{k}\sigma^{3},\qquad\omega_{k}=\frac{1}{2}e_{kl}\partial_{l}\rho$ (5.2)
@@ 14.4 Stochastic Quantization and Complex Langevin Dynamics
15 Catastrophe Theory and the Geometric Mass Spectrum
15.1 Twisted Boundary Conditions, Symmetrization, and Spin Projections
@@
Before solving the exact spectral equations, we justify the choice of the twisted geometry from the perspective of the semiclassical quark path integral. In Planar QCD, the effective Minkowski action involves a sum over quark trajectories weighted by the area of the confining minimal surface.
To extract the spectrum of states with macroscopic angular momentum $J$, we impose twisted boundary conditions. We require the quark positions at the ends of the string to undergo a physical spatial rotation by an angle $\phi$ over the period of the motion, with the angular momentum $J$ acting as the thermodynamic conjugate to this twist. The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime. The minimal surface spanning two twisted helical boundaries is the helicoid, as illustrated in Figure 14.
Figure 14: The helicoid spanned by a rotating $\overline{q}q$ pair connected by a rigid stick (string). This minimal surface bounded by a double helix, discovered by Meusnier in 1785, provides the macroscopic physical motivation for the twisted boundary conditions.
Thus, the helicoid is not merely an ansatz; it is the simplest geometric solution mandated by the kinematics of angular momentum in the semiclassical limit. To extract the physical spin states of the meson spectrum, we must couple the internal worldsheet geometry to this macroscopic angular momentum $J$ of the target space.
In our twistor string formulation, we project this twisted kinematics directly into the boundary twistor data. This is achieved by introducing the $SU(2)_{L}$ and $SU(2)_{R}$ projections, integrating over the target space rotation angle $\phi$ with the canonical spin weight:
$\mathcal{P}_{J}=\oint d\phi\exp\left(\mathrm{i}\phi J\right)$ (15.1)
When the macroscopic boundary of the string rotates by an angle $\phi$, the boundary twistors must geometrically twist. To enforce this, we introduce a fractional winding parameter $a$ corresponding to twisted boundary conditions for $\lambda(z)$.
Furthermore, because macroscopic QCD rigorously conserves parity, the physical effective action must be symmetrized over the Self-Dual (SD) and Anti-Self-Dual (ASD) minimal surfaces. In the twistor parameterization, this corresponds to activating the twisted boundary condition for $\lambda(z)$ while keeping the mirror sector $\mu(z)$ frozen to a constant basis spinor, and adding a mirror term with $\mu(z)$ to the exponent. The Hodge-duality (or anti-duality) explicitly breaks the 3D embedding constraint, forcing the chiral surface to genuinely span 4-dimensional Minkowski space.
To analytically generate the perfect resonance leading to the exact WKB saddle point, we must consider the limit of an infinitely large number of geometric windings $k\rightarrow\infty$. As the integration variable traverses the flat valley $k$ times, the accumulated action scales strictly linearly with $k$. Crucially, this means the transverse Hessian scales as $\mathcal{H}_{k}\propto k$. In the asymptotic resonance limit $k\rightarrow\infty$, the transverse restoring force becomes strictly infinite. The stochastic variance of the quantum fluctuations scales as $1/k\rightarrow 0$. The path integral dynamically localizes to the exact WKB saddle point, completely shielding the topological monodromy from quantum noise. These heuristic arguments will be made explicit and quantitative in this section, which will end with exact WKB quantization in a topological sector with only one pole of twistor field.
15.2 Geometric Quantization via Degenerate Lefschetz Thimbles
In the semiclassical (WKB) limit, the path integral over the twistor phase space is dominated by the saddle points $\mathcal{V}_{*}(E)$ of the complexified effective action. Physical bound states (mesons) correspond to macroscopic divergences in this amplitude. Mathematically, these occur at critical energies $E=E_{n}$ where the classical stability matrix (the Hessian $\mathcal{H}$) develops a null space.
Because the integration variables are coupled complex holomorphic twistors, the vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of exactly two real directions on the Lefschetz thimble (a strict corank-2 degeneracy). This complex holomorphic geometry natively upgrades the standard square-root branch cut into a strictly simple pole, precisely generating the discrete S-matrix mass poles of the meson spectrum without ad hoc approximations.
15.3 Catastrophes, flat valleys, and topological barriers
To dynamically generate an exact simple mass pole $\propto 1/(E_{n}-E)$, the steepest descent trajectory must correspond to a non-compact flat zero-mode that extends infinitely far in configuration space. This exact infinite flat valley is provided dynamically by the multi-valued monodromy of the holomorphic twistor fields as the boundary parameter winds around the enclosed twistor poles.
The integrity of this topological winding number is rigorously protected by an infinite topological action barrier. If a pole $z_{0}$ attempts to cross the unit circle from the inside ($|z_{0}|\to 1$), its conjugate mirror located at $1/\bar{z}_{0}$ must simultaneously approach the boundary from the outside. This creates a severe pinch singularity, strictly trapping the integration contour and locking the path integral into a conserved topological sector.
15.4 Poles inside the disk and the monodromies
For a saddle point with a single pole located strictly inside the unit disk ($|z_{0}|<1$), the total effective action accumulates strictly additively over multiple windings $k$. The Liouville bulk action simply duplicates, and the logarithmic boundary primitive picks up a strict topological phase shift at every cycle. Together, these terms yield an infinite geometric series $\sum\exp\left(\mathrm{i}k\Delta S_{eff}\right)$. The macroscopic divergence of this series generates the simple meson pole precisely when the exact Bohr-Sommerfeld quantization condition $\Delta S_{eff}=2\pi n$ is satisfied.
15.5 The Exact One-Pole Solution and Minkowski Transition
For the fundamental meson trajectory, we consider in detail the specific case of a single pole. To implement the twisted boundary condition for $\lambda(z)$ while freezing $\mu(z)$ to enforce the symmetrization over SD and ASD surfaces, we utilize the exact chiral 1-pole configuration with fractional power $a$, shifted to a pole position $z_{0}$:
$\lambda(z)=R(z-z_{0})^{-a}\begin{pmatrix}1\\ 0\end{pmatrix},\qquad\mu(z)=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ 1\end{pmatrix}$ (15.2)
As we trace the boundary loop $z\to z\exp\left(2\pi\mathrm{i}\right)$, the active twistor acquires the fractional monodromy $\lambda\to\exp\left(-2\pi\mathrm{i}a\right)\lambda$. Equating this to the physical spin-1/2 transformation $\lambda\to\exp\left(-\mathrm{i}\phi/2\right)\lambda$, we establish the exact geometric identification $\phi=4\pi a$. The angular momentum projection operator thereby injects the canonical phase $\exp\left(4\pi\mathrm{i}Ja\right)$.
To evaluate the bulk action properly, we must transition to the physical Minkowski metric on the symmetrized chiral surface, as established in D. We switch to the Minkowski metric on the surface parameterized by $z=\exp\left(\mathrm{i}\sigma_{M}+\tau_{E}\right)$, describing the Nambu-Goto metric $g_{ab}$ via the coordinate transformation $\exp\left(2\rho_{E}\right)d^{2}z=R^{2}\frac{d^{2}z}{(z\bar{z})^{a}}=R^{2}\mathrm{i}d\sigma_{M}d\tau_{E}$. Upon double Wick rotation and symmetrization (taking the parity-even arithmetic mean of the Left and Right chiral configurations), the conformal factor of the twisted metric dynamically restricts the causal domain. Using the modified radial coordinate $\tilde{\sigma}_{M}=(1-a)\sigma_{M}$, the physical speed-of-light boundary occurs exactly where this conformal metric vanishes, at $\tilde{\sigma}_{M}=\pi/2$. The proper Minkowski Area integral factorizes in the light-cone coordinates $z,\bar{z}$:
$\text{Area}(a)$ $=\frac{\mathrm{i}R^{2}}{2}\int_{z\overline{z}<1}dz\wedge d\overline{z}(z\overline{z})^{-a}$
$=\frac{-\mathrm{i}R^{2}}{2}\oint\frac{dz}{z}\int_{0}^{1}d(z\overline{z})(z\overline{z})^{-a}$
$=\frac{-\mathrm{i}R^{2}}{2(1-a)}\oint dzz^{-1}=\frac{\pi R^{2}}{1-a}$ (15.3)
Multiplying by the physical string tension $\sigma$, the macroscopic Area action evaluates exactly to:
$S_{M,Area}=\frac{\pi\sigma R^{2}}{1-a}$ (15.4)
Simultaneously, the Liouville action evaluates the kinetic gradient term. Crucially, while the target space metric stretches with the twist, the topological Liouville anomaly resides strictly on the rigidly bounded fundamental domain of the reference worldsheet, $\sigma_{M}\in[0,\pi/2]$:
$\int|\nabla\rho|^{2}d^{2}z$ $=a^{2}\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}d\sigma_{M}$
$=a^{2}(2\pi)\left(\frac{\pi}{2}\right)=\pi^{2}a^{2}$ (15.5)
Multiplying by the universal anomaly coefficient $\frac{1}{12\pi}$, the Minkowski kinetic action yields the strictly positive, rigid Lüscher shift:
$S_{M,kin}=+\frac{\pi a^{2}}{12}$ (15.6)
[…]
$\Delta X_{0}=\int_{0}^{2\pi}\sqrt{2}R\sin\left((1-a)\theta\right)d\theta=\frac{2\sqrt{2}R\sin^{2}(\pi a)}{1-a}$ (15.7)
We balance the string tension against the boundary energy by minimizing the combined classical action $S(R)=S_{M,Area}-E\Delta X_{0}$ with respect to the geometric scale $R$. Since the Area is purely quadratic $(AR^{2})$ and the boundary coupling is strictly linear $(BR)$, the minimization of this quadratic function trivially yields $B^{2}/(4A)$ by the virial theorem. This dynamically eliminates the dimensionful size $R$ in one step, leaving a purely intensive, remarkably pristine geometric boundary potential:
$S_{geom}(a)=-\frac{E^{2}}{\pi\sigma}F(a),\qquad\text{where}\qquad F(a)=\frac{2\sin^{4}(\pi a)}{1-a}$ (15.8)
This establishes the exact, scale-independent geometric potential that governs the twistor branch point degree $a$ at the WKB saddle point.
15.6 Quark Spin Monodromy, Trace Splitting, and the Unified Spectrum
The final component of the path integral is the path-ordered Dirac trace $T$ along the quark boundary. The global boundary holonomy matrix acting on the twistor space is $U=\text{diag}(\exp\left(-2\pi\mathrm{i}a\right),1)$.
1. The $\rho$ Trajectory (Vector Current)
With an extra $\gamma_{\nu}$ projector inserted as the vector quark current, the twistor dyadic projector $P\propto\lambda\mu^{\dagger}$ acts as a strict rank-1 projector, $\left(\begin{smallmatrix}1&1\\ 0&0\end{smallmatrix}\right)$. It annihilates the trivial constant mode of the un-twisted right-chiral twistor, leaving exactly:
$T_{\rho}\propto\text{tr}(UP)$ $=\text{tr}\begin{pmatrix}\exp\left(-2\pi\mathrm{i}a\right)&\exp\left(-2\pi\mathrm{i}a\right)\\ 0&0\end{pmatrix}$
$=\exp\left(-2\pi\mathrm{i}a\right)$ (15.9)
This corresponding pure topological phase factor $\exp\left(-2\pi\mathrm{i}a\right)$ adds directly to the rotation weight $\exp\left(4\pi\mathrm{i}Ja\right)$, shifting the angular momentum exactly by $\Delta J=+1/2$.
2. The $\pi$ and $K$ Trajectories (Pseudoscalar Current)
With a $\gamma_{5}$ inserted for the pseudoscalar and scalar trajectories, the trace evaluates the sum of the modes:
$T_{\pi,K}=\text{tr}(U)=1+\exp\left(-2\pi\mathrm{i}a\right)$ (15.10)
Because the path integral weight is linear in $T$, the amplitude rigorously splits into a sum of two independent path integrals: